# A magnetic avalanche as the central engine powering a solar flare

### v1

@@ 6 Conclusion
This unprecedented set of Solar Orbiter observations of an M7.7 flare showcases a comprehensive picture of the central engine of the preflare and impulsive phases of a solar flare in the form of a magnetic avalanche. Acceleration of non-thermal particles is closely related to the avalanche reconnection within the magnetic flux rope itself that contains the filament material. Energy transport signatures in the lower atmosphere are pervasively seen as raining plasma blobs over a long duration starting before the impulsive phase of the flare.
@@ Acknowledgements.
LPC gratefully acknowledges helpful discussions with Hugh Hudson, Hamish Reid, and Marc DeRosa. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement Nos. 10103984 – project ORIGIN; 101097844 – project WINSUN). DP gratefully acknowledges support from the Australian Research Council through the award of two Discovery Projects (DP210100709, DP230101240). DB, EK, LR, CV, ANZ thank the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA) under contract numbers 4000136424, 4000134474, 4000134088 and 4000143743. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. We thank the ESA SOC and MOC teams for their support. The EUI instrument was built by CSL, IAS, MPS, MSSL/UCL, PMOD/WRC, ROB, LCF/IO with funding from the Belgian Federal Science Policy Office (BELSPO/PRODEX PEA 4000134088, 4000106864 and 4000112292); the Centre National d’Etudes Spatiales (CNES); the UK Space Agency (UKSA); the Bundesministerium für Wirtschaft und Energie (BMWi) through the Deutsches Zentrum für Luft- und Raumfahrt (DLR); and the Swiss Space Office (SSO). The STIX instrument is an international collaboration between Switzerland, Poland, France, Czech Republic, Germany, Austria, Ireland, and Italy. The development of the SPICE instrument has been funded by ESA member states and ESA (contract no. SOL.S.ASTR.CON.00070). The German contribution to SPICE is funded by the Bundesministerium für Wirtschaft und Technologie through the Deutsches Zentrum für Luft- und Raumfahrt e.V. (DLR), grants no. 50 OT 1001/1201/1901. The Swiss hardware contribution was funded through PRODEX by the Swiss Space Office (SSO). The UK hardware contribution was funded by the UK Space Agency. The German contribution to SO/PHI is funded by the BMWi through DLR and by MPG central funds. The Spanish contribution is funded by AEI/MCIN/10.13039/501100011033/ and European Union “NextGenerationEU”/PRTR” (RTI2018-096886-C5, PID2021-125325OB-C5, PCI2022-135009-2, PCI2022-135029-2) and ERDF “A way of making Europe”; “Center of Excellence Severo Ochoa” awards to IAA-CSIC (SEV-2017-0709, CEX2021-001131-S). The French contribution is funded by CNES. SDO is the first mission to be launched for NASA’s Living With a Star (LWS) Program and the data supplied courtesy of the HMI and AIA consortia. We thank Fermi and GOES teams for making the data publicly available. This research has made use of NASA’s Astrophysics Data System Bibliographic Services. AIAPY open source software package v0.7.3 (Barnes et al. 2020, 2021), SSWIDL (Freeland & Handy 1998), sunpy version 6.0.4 (The SunPy Community et al. 2020), and mayavi (Ramachandran & Varoquaux 2011) are used for data analysis.References
@@ A.1 EUI/HRIEUV data and time-distance diagrams
The high-resolution imager of the Extreme Ultraviolet Imager (EUI/HRIEUV) instrument (Rochus et al. 2020) has a detector of 3072 3072 pixels of which the central 2048 2048 pixels are illuminated and read-out. It observed the flare with a high cadence of 2 s at 1.65 s exposure time, and an image scale of 0.492 arcseconds (′′) pixel-1, where 1″212 km on the Sun at 0.2926 au (Fig. 5). Therefore, the spatial resolution of the HRIEUV (core of the point spread function of the instrument has a full width at half maximum of 2 pixels) was about 209 km at the time of observations. The HRIEUV filter, centered around 174 Å, has a dominant contribution from the EUV emission lines of Fe ix (at 171.1 Å) and Fe x (at 174.5 Å and 177.2 Å). The filter has its response peak around 1 MK. Accordingly, the event is one of the highest resolution coronal flares ever recorded in the EUV. We made use of the level-2 HRIEUV data in our analysis (Kraaikamp et al. 2023). The spacecraft jitter in the EUV images was removed using a cross-correlation technique detailed in Chitta et al. (2022, 2023).
In Figs. 2 and 3, we show the time-distance diagrams along selected artificially placed slits on the EUI images (Fig. 6). In each case, we considered pixels perpendicular to the slit direction, averaged emission over this width and then stacked the resulting mean intensity along the slit as a function of time.
@@ A.4 STIX data analysis
The Spectrometer Telescope for Imaging X-rays (STIX; Krucker et al. 2020; Xiao et al. 2023) on board Solar Orbiter detects hard X-rays in the energy range 4–150 keV, with imaging and spectroscopic capabilities. STIX captured the initial and impulsive phases of the flare at a high cadence of 0.5 s. The decay phase after 2024 October 01 UT 00:25:54 was recorded at 2 s cadence (Fig. 5). We computed STIX count rates in 5 different energy bins, ranges of which are indicated by the colored labels in Fig. 1.
@@
We constructed images of the STIX hard X-ray sources to investigate their spatial localization with respect to the coronal dynamics recorded by the HRIEUV instrument. We used the standard STIX procedures and routines available in SolarSoft for this purpose. In particular, for the X-ray image reconstruction we employed a maximum entropy method (Massa et al. 2020). To account for the absolute pointing uncertainties of the HRIEUV and STIX instruments in addition to their relative pointing uncertainties, we supplied the helioprojective cartesian coordinates of the flare in Solar Orbiter frame as (3057.9″,-930.7″) to the STIX X-ray image reconstruction procedures. We constructed six STIX images, with pixel size of 4″, over a field of view of pixels, covering different phases of the flare in specific energy bins (e.g., contours shown in Fig. 1 and the associated online movie). In each of these phases, we integrated STIX data over different intervals depending on the signal. The integration time is indicated in the movie and explained in the caption of Fig. 1. We further confirmed that the hard X-ray sources presented in the movie remain robust even if we increase the spatial resolution by decreasing the pixel size to 2″, or even 1″.
@@
Furthermore, we employed the Object Spectral Executive (Schwartz et al. 2002) to investigate the characteristics of STIX hard X-ray sources at the peak of impulsive phase of the flare (a 45 s interval starting from 2024-Sep-30 UT 23:47:03, when the STIX signal in the 25–50 keV and 50–76 keV energy bands exhibits a clear intensity peak; see contours in Fig. 1). We first subtracted a pre-flare background from the spectrum (background interval of 180 s starting from 2024-Sep-30 UT 23:36:00) and fitted the resulting data with an isothermal model assuming coronal abundances (Feldman 1992) along with a cold thick target non-thermal emission model (Holman et al. 2011). This non-thermal emission is theorized to arise from an electron population above a low-energy cutoff , having a negative power-law distribution (with index ). The resulting plasma temperature, emission measure, and along with the model fits are shown in Fig. 9. Immediately after this impulsive phase, STIX light curves, particularly in the 25–50 keV energy range exhibited oscillatory behavior. Based on the image reconstruction, we found that the signal originates from coronal altitudes at the apexes of flaring loops. We fitted the STIX spectrum, integrated over an interval of 78 s starting from 2024-Sep-30 UT 23:47:55, with a thermal and thin-target model (Brown 1971).
@@ A.5 Solar Dynamics Observatory data and magnetic connectivity
To generally understand this coronal large-scale magnetic connectivity, we employed magnetic field extrapolations. But because the active region is away from the disk center and given the complex surface magnetic field distribution over larger-scales of over 200 Mm, the curvature of the Sun cannot be ignored. This means that local extrapolations covering only the active region may not be adequate. To mitigate this, we have considered employing global extrapolations of the surface magnetic field. To this end, we first retrieved the surface field closer in time to the flare hard X-ray peak from the Solarsoft Potential Field Source Surface module (Schrijver & De Rosa 2003). This surface field is constructed at a scale of 0.5° pixel-1 in longitude () and latitude () directions (upper panel of Fig. 12). We then used a finite difference magnetic field extrapolation technique to construct the coronal field (Tóth et al. 2011). To this end, we considered 172 grid points in the radial direction, with the the upper boundary (source surface) placed at (where solar radius, Mm).
@@
Analysis of the magnetic field structure shown in Figures 12 and 13 was performed using Universal Fieldline Tracer (UFiT; https://github.com/Valentin-Aslanyan/UFiT; Aslanyan et al. (2024)). The signed-log (slog) of the magnetic squashing factor, (Titov 2007; Scott et al. 2017), defined to be positive for closed field lines and negative for open field lines, is presented in the lower panel of Fig. 12. High reveals the presence of strong gradients in the field line mapping, that correspond to either true separatrix surfaces or quasi-separatrix layers. We then identified magnetic nulls (regions prone to magnetic reconnection) using a trilinear null-finding method (Chiti et al. 2020). We mark the locations of three null points in the inset of the lower panel of Figure 12. These three nulls are located at altitudes between 16 and 30 Mm above the photosphere in the present extrapolation (their existence appears to be relatively robust to changing the time or resolution of the magnetogram employed, though the details of their positions vary). We then traced magnetic field lines passing close to the three nulls (Fig. 13). Their footpoints align with many of the observed remote brightenings. Also visible is the footpoint of the flaring filament, located within the lower arcade below this structure. The trace of these arcades can be seen as sharp high- lines in the inset of Figure 12, where the “cross-hairs” mark the approximate location of the filament footpoint. Although the potential field extrapolations are rather simplistic, there is a good observational correspondence of the traced field lines. In this scenario, the remote flare ribbons would have been caused by energy transfer from the flare site along the field lines from the null points and/or separators connecting them to one another where the reconnection was likely occurring (Pontin & Priest 2022). The flux rope breaks through this surface, enabled by the reconnection associated with the jets seen in Figure 3, causing the flux rope to become ‘disconnected’ from the Sun at its southern end (see also Karpen et al. 2024). This disconnection process triggers rapid dynamics within the erupting flux rope.
@@ A.6 Fermi and GOES data
As the flare happened on the side of the solar disk facing Earth, the Fermi Gamma-ray Burst Monitor (Fermi/GBM; Meegan et al. 2009) could detect the hard X-ray emission with its Na i detectors, which are sensitive to energies above a few keV up to about 1 MeV. We computed Fermi/GBM count rates in the same energy bins that we used for STIX data by integrating the signal from the three most sunward facing detectors and subtracting from it the signal from the three least sunward facing detectors. The Fermi/GBM light curve in the 25–50 keV energy bin is shown in Fig. 3E. In Fig. 5 we plot the soft X-ray flux measured by the 1–8 Å filter on the Geostationary Operational Environmental Satellite (GOES). Both the Fermi and GOES data are publicly available and are retrieved using the standard routines available via solarsoft.
@@ Appendix B Additional discussion on HXR sources
In Fig. 1 we observed that the highest energy STIX contour in the 50–76 keV range is spatially offset from the other components. This offset is robust with image reconstructions at different spatial resolutions. In particular, the 50–76 keV source region is at least a few Mm away from the other sources. This would imply that the higher energy particles have precipitated to produce the bremsstrahlung emission from a deeper atmospheric layer than the other sources. At the same time, we identified compact enhancements in the photosphere, particularly in the line core images sampled by the PHI instrument. Moreover, the 25–50 keV time series of STIX exhibited an extended secondary enhancement, as seen in the count rates during the post-impulsive phase, for about 5–10 minutes. There is a potential signature of this also in the PHI observations (maroon light curve in Fig. 8) in a compact region at the footpoint of the filament. There are also indications of ribbon line bright front propagation over the penumbral region of the sunspot (Movie S7). These enhancements are likely signatures of strong photospheric heating that could be caused by accelerated particles depositing their energy in the deeper atmosphere (Granovsky et al. 2024).
@@
Appendix C Captions for Movies

### v2

@@ 6 Conclusion
This unprecedented set of Solar Orbiter observations of an M7.7 flare showcases a comprehensive picture of the central engine of the pre-flare and impulsive phases of a solar flare in the form of a magnetic avalanche. The acceleration of nonthermal particles is closely related to the avalanche reconnection within the magnetic flux rope itself that contains the filament material. Energy transport signatures in the lower atmosphere are pervasively seen as raining plasma blobs over a long duration starting before the impulsive phase of the flare.
@@ Acknowledgements.
The authors thank the anonymous referee for constructive comments that helped improve the presentation of the manuscript. LPC gratefully acknowledges helpful discussions with Hugh Hudson, Hamish Reid, and Marc DeRosa. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement Nos. 10103984 – project ORIGIN; 101097844 – project WINSUN). We also acknowledge funding received under the Horizon Europe program of the European Union (grant agreement no. 101131534 – project DynaSun). DP gratefully acknowledges support from the Australian Research Council through the award of two Discovery Projects (DP210100709, DP230101240). DB, EK, LR, CV, ANZ thank the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Program of the European Space Agency (ESA) under contract numbers 4000136424, 4000134474, 4000134088 and 4000143743. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. We thank the ESA SOC and MOC teams for their support. The EUI instrument was built by CSL, IAS, MPS, MSSL/UCL, PMOD/WRC, ROB, LCF/IO with funding from the Belgian Federal Science Policy Office (BELSPO/PRODEX PEA 4000134088, 4000106864 and 4000112292); the Centre National d’Etudes Spatiales (CNES); the UK Space Agency (UKSA); the Bundesministerium für Wirtschaft und Energie (BMWi) through the Deutsches Zentrum für Luft- und Raumfahrt (DLR); and the Swiss Space Office (SSO). The STIX instrument is an international collaboration between Switzerland, Poland, France, Czech Republic, Germany, Austria, Ireland, and Italy. The development of the SPICE instrument has been funded by ESA member states and ESA (contract no. SOL.S.ASTR.CON.00070). The German contribution to SPICE is funded by the Bundesministerium für Wirtschaft und Technologie through the Deutsches Zentrum für Luft- und Raumfahrt e.V. (DLR), grants no. 50 OT 1001/1201/1901. The Swiss hardware contribution was funded through PRODEX by the Swiss Space Office (SSO). The UK hardware contribution was funded by the UK Space Agency. The German contribution to SO/PHI is funded by the BMWi through DLR and by MPG central funds. The Spanish contribution is funded by AEI/MCIN/10.13039/501100011033/ and European Union “NextGenerationEU”/PRTR” (RTI2018-096886-C5, PID2021-125325OB-C5, CNS2023-144723) and ERDF “A way of making Europe”; “Center of Excellence Severo Ochoa” awards to IAA-CSIC (SEV-2017-0709, CEX2021-001131-S). The French contribution is funded by CNES. SDO is the first mission to be launched for NASA’s Living With a Star (LWS) Program and the data supplied courtesy of the HMI and AIA consortia. We thank Fermi and GOES teams for making the data publicly available. This research has made use of NASA’s Astrophysics Data System Bibliographic Services. AIAPY open source software package v0.7.3 (Barnes et al. 2020, 2021), SSWIDL (Freeland & Handy 1998), sunpy version 6.0.4 (The SunPy Community et al. 2020), and mayavi (Ramachandran & Varoquaux 2011) are used for data analysis.References
@@
- Cheng et al. (2023) Cheng, X., Priest, E. R., Li, H. T., et al. 2023, Nature Communications, 14, 2107
- Cheung et al. (2019) Cheung, M. C. M., Rempel, M., Chintzoglou, G., et al. 2019, Nature Astronomy, 3, 160
@@ A.1 EUI/HRIEUV data and time-distance diagrams
The high-resolution imager of the Extreme Ultraviolet Imager (EUI/HRIEUV) instrument (Rochus et al. 2020) has a detector of 3072 3072 pixels of which the central 2048 2048 pixels are illuminated and read-out. It observed the flare with a high cadence of 2 s at 1.65 s exposure time, and an image scale of 0.492 arcseconds (′′) pixel-1, where 1″212 km on the Sun at 0.2926 au (Fig. 6). Therefore, the spatial resolution of the HRIEUV (core of the point spread function of the instrument has a full width at half maximum of 2 pixels) was about 209 km at the time of observations. The HRIEUV filter, centered around 174 Å, has a dominant contribution from the EUV emission lines of Fe ix (at 171.1 Å) and Fe x (at 174.5 Å and 177.2 Å). The filter has its response peak around 1 MK. Accordingly, the event is one of the highest-resolution coronal flares ever recorded in the EUV. We made use of the level-2 HRIEUV data in our analysis (Kraaikamp et al. 2023). The spacecraft jitter in the EUV images was removed using a cross-correlation technique detailed in Chitta et al. (2022, 2023).
In Figs. 3 and 4, we show the time-distance diagrams along selected artificially placed slits on the EUI images (Fig. 2). In each case, we considered pixels perpendicular to the slit direction, averaged emission over this width and then stacked the resulting mean intensity along the slit as a function of time.
@@ A.4 STIX data analysis
The Spectrometer Telescope for Imaging X-rays (STIX; Krucker et al. 2020; Xiao et al. 2023) on board Solar Orbiter detects HXRs in the energy range 4–150 keV, with imaging and spectroscopic capabilities. STIX captured the initial and impulsive phases of the flare at a high cadence of 0.5 s. The decay phase after 2024 October 01 UT 00:25:54 was recorded at 2 s cadence (Fig. 6). We computed STIX count rates in 5 different energy bins, ranges of which are indicated by the colored labels in Fig. 1.
@@
We constructed images of the STIX HXR sources to investigate their spatial localization with respect to the coronal dynamics recorded by the HRIEUV instrument. We used the standard STIX procedures and routines available in SolarSoft for this purpose. In particular, for the X-ray image reconstruction we employed a maximum entropy method (Massa et al. 2020). To account for the absolute pointing uncertainties of the HRIEUV and STIX instruments in addition to their relative pointing uncertainties, we supplied the helioprojective cartesian coordinates of the flare in Solar Orbiter frame as (3057.9″,-930.7″) to the STIX X-ray image reconstruction procedures. We constructed six STIX images, with pixel size of 4″, over a field of view of pixels, covering different phases of the flare in specific energy bins (e.g., contours shown in Fig. 1 and the associated online movie). In each of these phases, we integrated STIX data over different intervals depending on the signal. The integration time is indicated in the movie and explained in the caption of Fig. 1. We further confirmed that the HXR sources presented in the movie remain robust even if we increase the spatial resolution by decreasing the pixel size to 2″, or even 1″.
@@
Furthermore, we employed the Object Spectral Executive (Schwartz et al. 2002) to investigate the characteristics of STIX HXR sources at the peak of impulsive phase of the flare (a 45 s interval starting from 2024-Sep-30 UT 23:47:03, when the STIX signal in the 25–50 keV and 50–76 keV energy bands exhibits a clear intensity peak; see contours in Fig. 1). We first subtracted a pre-flare background from the spectrum (background interval of 180 s starting from 2024-Sep-30 UT 23:36:00) and fitted the resulting data with an isothermal model assuming coronal abundances (Feldman 1992) along with a cold thick target nonthermal emission model (Holman et al. 2011). This nonthermal emission is theorized to arise from an electron population above a low-energy cutoff , having a negative power-law distribution (with index ). The resulting plasma temperature, emission measure, and along with the model fits are shown in Fig. 9. Immediately after this impulsive phase, STIX light curves, particularly in the 25–50 keV energy range exhibited oscillatory behavior. Based on the image reconstruction, we found that the signal originates from coronal altitudes at the apexes of flaring loops. We fit the STIX spectrum, integrated over an interval of 78 s starting from 2024-Sep-30 UT 23:47:55, with a thermal and thin-target model (Brown 1971).
@@ A.5 Solar Dynamics Observatory data and magnetic connectivity
To generally understand this coronal large-scale magnetic connectivity, we employed magnetic field extrapolations. But because the active region is away from the disk center and given the complex surface magnetic field distribution over larger-scales of over 200 Mm, the curvature of the Sun cannot be ignored. This means that local extrapolations covering only the active region may not be adequate. To mitigate this, we have considered employing global extrapolations of the surface magnetic field. To this end, we first retrieved the surface field closer in time to the flare HXR peak from the Solarsoft Potential Field Source Surface module (Schrijver & De Rosa 2003). This surface field is constructed at a scale of 0.5° pixel-1 in longitude () and latitude () directions (upper panel of Fig. 12). We then used a finite difference magnetic field extrapolation technique to construct the coronal field (Tóth et al. 2011). To this end, we considered 172 grid points in the radial direction, with the the upper boundary (source surface) placed at (where solar radius, Mm).
Analysis of the magnetic field structure shown in Figures 12 and 13 was performed using Universal Fieldline Tracer (UFiT; https://github.com/Valentin-Aslanyan/UFiT; Aslanyan et al. (2024)). The signed-log (slog) of the magnetic squashing factor, (Titov 2007; Scott et al. 2017), defined to be positive for closed field lines and negative for open field lines, is presented in the lower panel of Fig. 12. High reveals the presence of strong gradients in the field line mapping, that correspond to either true separatrix surfaces or quasi-separatrix layers. We then identified magnetic nulls (regions prone to magnetic reconnection) using a trilinear null-finding method (Chiti et al. 2020).
@@
We mark the locations of three null points in the inset of the lower panel of Figure 12. These three nulls are located at altitudes between 16 and 30 Mm above the photosphere in the present extrapolation (their existence appears to be relatively robust to changing the time or resolution of the magnetogram employed, though the details of their positions vary). We then traced magnetic field lines passing close to the three nulls (Fig. 13). Their footpoints align with many of the observed remote brightenings. Also visible is the footpoint of the flaring filament, located within the lower arcade below this structure. The trace of these arcades can be seen as sharp high- lines in the inset of Figure 12, where the “cross-hairs” mark the approximate location of the filament footpoint. Although the potential field extrapolations are rather simplistic, there is a good observational correspondence of the traced field lines. In this scenario, the remote flare ribbons would have been caused by energy transfer from the flare site along the field lines from the null points and/or separators connecting them to one another where the reconnection was likely occurring (Pontin & Priest 2022). The flux rope breaks through this surface, enabled by the reconnection associated with the jets seen in Figure 4, causing the flux rope to become ”disconnected” from the Sun at its southern end (see also Karpen et al. 2024). This disconnection process triggers rapid dynamics within the erupting flux rope.
@@ A.6 Fermi and GOES data
As the flare happened on the side of the solar disk facing Earth, the Fermi Gamma-ray Burst Monitor (Fermi/GBM; Meegan et al. 2009) could detect the HXR emission with its Na i detectors, which are sensitive to energies above a few keV up to about 1 MeV. We computed Fermi/GBM count rates in the same energy bins that we used for STIX data by integrating the signal from the three most sunward facing detectors and subtracting from it the signal from the three least sunward facing detectors. The Fermi/GBM light curve in the 25–50 keV energy bin is shown in Fig. 4E. In Fig. 6 we plot the soft X-ray flux measured by the 1–8 Å filter on the Geostationary Operational Environmental Satellite (GOES). Both the Fermi and GOES data are publicly available and are retrieved using the standard routines available via solarsoft.
@@ Appendix B Additional discussion on HXR sources
In Fig. 1 we observed that the highest-energy STIX contour in the 50–76 keV range is spatially offset from the other components. This offset is robust with image reconstructions at different spatial resolutions. In particular, the 50–76 keV source region is at least a few Mm away from the other sources. This would imply that the higher-energy particles have precipitated to produce the bremsstrahlung emission from a deeper atmospheric layer than the other sources. At the same time, we identified compact enhancements in the photosphere, particularly in the line core images sampled by the PHI instrument. Moreover, the 25–50 keV time series of STIX exhibited an extended secondary enhancement, as seen in the count rates during the post-impulsive phase, for about 5–10 minutes. There is a potential signature of this also in the PHI observations (maroon light curve in Fig. 8) in a compact region at the footpoint of the filament. There are also indications of ribbon line bright front propagation over the penumbral region of the sunspot (Movie S7). These enhancements are likely signatures of strong photospheric heating that could be caused by accelerated particles depositing their energy in the deeper atmosphere (Granovsky et al. 2024).
@@
Appendix C Discussion on the MHD models of a magnetic avalanche
Magnetic avalanche reconnection can be considered a truly universal phenomenon. It is not only invoked to explain the power-law distribution of solar and stellar flares, and nanoflare heating of solar coronal loops, but also reconnection cascades in the heliospheric current sheet (Eriksson et al. 2022) and even magnetic flux eruptions in Kerr black hole simulations (Nalewajko et al. 2024).
In an MHD avalanche scenario in the solar context, a small perturbation to a marginally stable magnetic strand can trigger reconnection, disturbances from which propagates in space and time to neighboring magnetic strands that release stored energy by further reconnection. This gives rise to a catastrophic release of magnetic energy on short timescales. In our observations, we catch this phenomenon during the pre-flare and impulsive phases of the flare using coronal observations whose high spatiotemporal resolution is unprecedented.
MHD avalanches in solar coronal loops have to date been modeled in a simplified magnetic configuration, that is a straightened flux tube (Hood et al. 2016; Reid et al. 2023). These models are primarily used to investigate the nanoflare theory of coronal heating. Nanoflares possess only a millionth to billionth of the energy content of the type of M-class flare we are investigating here. As such, the model setup of a straightened loop does not allow for the buildup of the large amounts of energy necessary to produce a major flare event. Realistic magnetic configurations with sunspots subjected to large-scale shear, flux emergence and cancellation are required for this.
To this end, some of the best simulations of flares to date are produced by MURaM, a radiative MHD code (Cheung et al. 2019; Rempel et al. 2023). These simulations have a grid spacing of about 192 km in the horizontal direction and produce an overall realistic-looking flare. Although those simulations were not analyzed to investigate the onset of a magnetic avalanche, Figure 4 and the associated movie in Rempel et al. (2023) already give tantalizing hints of numerous small-scale reconnection events in the erupting flux rope in the pre-flare phase, similar to our observations (Fig. 3 in the main text). But the simulations need to be carefully analyzed at much higher spatial resolution to fully understand the coronal evolution on small spatial scales. Also, because they lack the treatment of nonthermal particles, the simulations do not yet reveal all the small-scale structuring of flare ribbons that we were able to identify with our EUI observations. On the one hand, there is already a good indication that the existing large-scale MHD models hint at features similar to our observations. On the other hand, a lot of progress still needs to be made in understanding of the onset of magnetic reconnection and particle acceleration in simulations – and the MHD framework alone will not yield this. Simulations at even higher grid spacing coupled with at least approximations of kinetic effects are needed to reproduce the spatial and temporal evolution of a flare that we are able to unravel with our analysis.
Appendix D Captions for movies